Status Report

NASA Space Station On-Orbit Status 10 November 2004

By SpaceRef Editor
November 10, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 10 November 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

First thing after breakfast, FE Sharipov terminated the bake-out cycle on the BMP micropurification system’s channel 2, moding it back to Purify.  This restored both filter beds to Purify mode.  [The regeneration of the air purifier filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]

Afterwards (~3:15am EST), Sharipov powered up the Elektron O2 generator, which shares the vacuum vent valve (in its case for H2) with the BMP.

Later in the day, Salizhan also performed the periodic replenishing of the Elektron’s condensate water supply for electrolysis, his third, filling the KOV thermal loop’s EDV container with purified (deionized) water from the BKO multifiltration/purification column unit.

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

CDR/SO Leroy Chiao printed out and reviewed the troubleshooting procedures for the failed EMU #3005, then began the scheduled repair activity on the spacesuit, viz., to install a new impeller/rotor in its water pump.  However, the work could not be completed due to a missing piece of hardware (a small laminated shim).  The suit was “buttoned up” for the time being, and ground specialists are developing new options.  If the shim cannot be located, another unit will be manifested on the next Progress (16P, to be launched 12/23).

After smoke detector (SD) auto-response for the FGB was inhibited by the onboard automated command sequencer (SPP), Sharipov swapped out two IDZ-2 SDs in Zarya (#5 & #10), first demating, then reconnecting their BKS cables.  Afterwards, the FGB fire detection system was reactivated and tested, before it was moded to Monitoring.

Chiao conducted Part 2 of the current BCAT-3 (Binary Colloid Alloy Test 3) science photography session, a repeat of an earlier run on 11/2.   [Unexpected results were seen by the PI (Principal Investigator) on inspecting the results of the earlier runs for Samples 1 & 2, which appear to have been insufficiently mixed.  After setting up the SGSM (slow growth sample module) in the Lab MWA at EXPRESS Rack 5 (ER5) yesterday and imaging Sample 1, the Science Officer today conducted the video-recorded activity of photographing Sample 1 at the 100,000 sec interval, followed by homogenization (with a bar magnet) and photography of Sample 2.  The images captured during the photo sessions are kept on separate PCMCIA memory cards until their successful archiving on the ground is confirmed.]

The CDR deployed two passive FMK (formaldehyde monitoring kit) sampling assemblies in the Lab (below CEVIS) and SM (most forward handrail), to catch any atmospheric formaldehyde on a collector substrate for subsequent analysis.  (Last time done: 10/13).

Additional cabin air sampling with the Russian AK-1M sampler was set up by the FE in the Service Module (SM) and in FGB/Zarya.  In Zvezda, Salizhan also checked for CO (carbon monoxide) with the IPD-CO Draeger tube samplers for report to TsUP/Moscow.  (Last time done:  10/14).] 

Using the WMK (water monitoring kit), Chiao conducted the visual T+2d analysis of the Week 3 water samples, collected on 11/8 from the SRV-K hot tap and the EDV container of the water supply system (SVO-ZV).  Part of the activity was another microbial analysis for inflight coliform bacteria (Escherichia coli) detection on samples collected on 10/8 (T+44h).  He then entered the microbiological data in the medical equipment computer.   [The analyses use incubated MCDs (microbial capture devices), SSK (surface sample kit) slides, and MAS (microbial air sampler) Petri dishes.  If Leroy’s analysis showed colony growth above specified limits, he was to take digital documentary images.  The used MCDs were discarded.]

Sharipov took photographs in the FGB of and behind specific stowage panels, to help the ground assess storage space availability.  Today’s imaging, to be continued tomorrow, focused on FGB panels 402, 403, 406, 407 and 410.

Salizhan spent time in the Soyuz-215/9S crew return vehicle to investigate an issue with “Neptune ME” control panel operation reported during the powered flight phase of 9S on 10/15 (Day 2).  Today’s task, which was supported by tagup with ground specialists on Daily Orbits 15 & 16, called for a performance check of the Neptune panel.   [Basically, during the ascent of the Soyuz spacecraft, some bad data showed up on its displays (inaccurate pressure readings).  Today’s tests by Sharipov indicate that this probably was a configuration issue between two displays that has been characterized and is now understood.  NASA engineers expect to see a procedural change forthcoming based on the test results.]

The CDR completed the regular weekly filling-out of the FFQ (food frequency questionnaire), his third, which keeps a log of his nutritional intake over time on special MEC (medical equipment computer) software.  The crew’s food ration audit yesterday is currently used at TsUP for putting together the food allocations to be launched on Progress 16P.   [The FFQ provides for recording consumed amounts during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins.]

The FE conducted the routine inspection of the SM’s SOZh environment control & life support system.

The crew performed their regular daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser, CEVIS bike ergometer, and VELO cycle with bungee cord load trainer.  Afterwards, Chiao downloaded TVIS and RED exercise data files to the MEC (medical equipment computer), currently six times a week to support early review on the ground.

From the discretionary Russian task list, Salizhan performed the periodic inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment that researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.

After yesterday’s work on the Russian ASN-M satellite navigation system by the FE to repair ASN feeder circuitry, TsUP is planning a test of the system for 11/19 after the transition to LVLH flight attitude.

As final step in the recent 4B2 battery reconditioning, MCC-Houston is today conducting a 24-hr. battery capacity test with the BCDU (battery charge/discharge unit) 4B2, which started at 9:00am EST and is running through 9:00am tomorrow.  Afterwards, 4B2 will be brought back on line.

The #2 SIGI (space integrated GPS/inertial navigation system) software upgrade with R2 firmware (GPS2) that was aborted on 11/4 after loading only the first of three loads will be attempted again tomorrow.  Currently, both SIGI-1 and Russian sources remain available.  SIGI-1 firmware load is scheduled for 11/16.  [Cause of the abort was a misconfiguration of the 760XDs comm ports due to different BIOS versions of the onboard SSCs (station support computers) and the ground SSC used to develop the procedure.  The latter’s BIOS has meanwhile been updated and tested.]

At ~9:10am, TsUP used the S-band to uplink a software modification for the TsVM Central Computer in the SM.  The patch is intended to prevent erratic FGB smoke detector telemetry from triggering a false smoke or fire event.

Starting at 3:00am EST this morning, MCC-Houston and HSG/Moscow repeated the 6-hr. BCC (Backup Control Center) checkout dry run that was only partially successful during the first try on 11/8.  Actual checkout operations are scheduled to start tomorrow 4:45am EST.

The Russian #1 air conditioner (SKV-1) continues to be down after its deactivation on 11/6 when it had failed to produce any condensate water in three days of operation.  Troubleshooting at TsUP continues.  SKV-2 is running nominally.

After yesterday’s dramatic drop in ISS altitude due to an energetic solar particle event with >100MeV protons (measured on the GOES 11 satellite), resulting in a massive spike in geomagnetic activity and in atmospheric drag at orbital altitude, solar activity appears to have leveled off and come down, with reduced orbital decay.  The crew was not in any danger.   [Yesterday’s drop in mean flight altitude reached ~440 m in 24 hrs, as compared to previous ~160 m/day.  This morning, the 24-hr. drop was ~200 m.  Next week’s planned reboost (11/17) will make up for the loss.]

Today’s CEO (crew earth observations) photo targets, limited currently by non-availability of the science window during 11/8-11/12 due to SSRMS operations, were Internal Waves, Timor Sea & Arafura Sea (weather was predicted to be clear to the NNW of Australia for internal wave photography.  Looking to the left of track and slightly behind for the sunglint point), Internal waves, Patagonian Shelf (a predicted clear weather window provided an opportunity for internal wave photography.  Looking to the left of track along the coastline for the sunglint point), and Lower Amazon River Basin, S. America (this overpass provided an opportunity for oblique photography of the northwestern portion of the Amazon River delta.  Photography of the estuary shoreline and islands to the left of track is important to detect changes in hydrologic and sedimentary processes).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events: 

  • Current station attitude of sun-pointing XPOP (x-axis perpendicular to orbit plane) will be maintained until 11/16;
  • Station reboost will be on 11/17 (backup opportunity 11/18);
  • Soyuz relocation is scheduled for 11/29 (no earlier than), preceded by a hot-fire test of its thrusters on 11/27;
  • Progress 15P is scheduled for undocking and destructive reentry on 12/22;
  • Progress 16P launch: 12/23; docking: 12/25 (GMT).
  •  U.S. & Russian Segment Status  (as of yesterday, 5:03pm EST)

  Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is Off (turned on this morning).  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is On.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is On.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off (failed).  SKV-2 is On (SM panel mods completed 4/8; SKV-2 activation failed 4/20; new replaceable condensate removal line installed on 9/9).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 755; temperature (deg C) — 26.3; ppO2 (mmHg) — 159.6; ppCO2 (mmHg) — 3.8.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 23.0.
  • Node:  Pressure (mmHg) — 751.5; temperature (deg C) — 24.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 752.9; temperature (deg C) — 22.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 753.1; temperature (deg C) — 26.3; ppO2 (mmHg) — 164.1; ppCO2 (mmHg) – 3.9.

(n/a = data not available.  U.S. ppO2 & ppCO2 readings from MCA)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack mode (suntracking).  Swiveled to 0 deg today for drag reduction.
  • SM batteries:  Battery #5 is in “Cycle” mode; all other batteries (7) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #6 is off line; battery #1 is in “Cycle” mode; all other batteries (4) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

  Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is failed, C&C-2 is prime, and C&C-3 is backup.
  • GNC-1 MDM (vers. R4) is prime; GNC-2 (vers. R4) is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

  Propulsion System (PS):

  • Total propellant load available: 3951 kg (8710 lb) as of 11/4/04;  [SM(552) + FGB(3399) + Progress M(0)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

  Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source US SIGI-1 (GPS)
  • Attitude source US String 1
  • Angular rate source — RGA-1

  Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = sun-fixed [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist) until 11/17.

Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:06am EST [= epoch]):

  • Mean altitude — 356.9 km
  • Apogee height — 362.4 km
  • Perigee height — 351.4 km
  • Period — 91.68 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0008125
  • Solar Beta Angle — 32.2 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.71
  • Mean altitude loss in last 24 hours — 200 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 34129

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.